OpenStudy (anonymous):

If dy/dx = 5x^2y and y(0)=6, find an equation for y in terms of x

myininaya (myininaya):

fun we have a function of y times a function of x=y' so we can use separation of variables

myininaya (myininaya):

so we have dy=5x^2y dx => dy/y=5x^2 dx => integrate both sides: lny=5x^3/3+C

myininaya (myininaya):

Now if you want to solve y just base e both sides giving you y=e^(5x^3/3+C)

myininaya (myininaya):

Ok but we can find C since you are given the initial condition y(0)=6

myininaya (myininaya):

y(0)=e^(0+C)=e^C=6 => so we have e^C=6 or C=ln6

myininaya (myininaya):

y=e^(5x^3/3+ln6)

myininaya (myininaya):

Now wasn't that fun? :)